Solve each system by addition:determine if the equation is independent, dependent or inconsistent
x-y=3
-6x+6y=17

Answer by stanbon(75887)

x-y=3
-6x+6y=17
---------------
Solve each for "y":
y = x -3
y = x + (17/6)
-------------------
Since both have a slope of "1" and unequal intercepts the lines are parallel.
The system in inconsistent.
======================
Cheers,
Stan H.